What is the kinetic energy of a car weighing 3000 pounds traveling at 55 miles per hour
Answer #1

Kinetic energy is given by= (1/2)(mass of object)(speed of object)2

in given unit's

Kinetic energy=  (1/2)(3000pounds)(55 miles/hr)2

= 4537500 pounds miles/hr

In SI units

M= 3000 pounds= 1360.777 kg

V= 55 miles/hour= 88513.9/3600= 24.5871 m/s

Kinetic energy= (1/2)(1360.777)(24.5871)2

= 411315.3488 J